Ask your own question, for FREE!
Mathematics 11 Online
OpenStudy (hockeychick23):

Tammy is playing a game where she is trying to roll a three with a standard die. If she gets a three in any of her first 4 rolls, she wins; otherwise she loses. What is the probability that Tammy wins the game?

OpenStudy (mathmate):

Hint: It's easier to calculate P(lose) = probability that she loses. Then P(win)=1-P(lose).

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
8 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
11 hours ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Arizona: does anyone know the percentage rate of 30,000 and 40,000
1 week ago 0 Replies 0 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!